A student has 540.0 mL of a 0.1115 M aqueous solution of CoCl2 to use in an experiment. He accidentally leaves the container uncovered and comes back the next week to find only a solid residue. The mass of the residue is 9.987 g. Determine the chemical formula of this residue.

If the substance is a hydrate, use a period instead of a dot in its formula.
1
Expert's answer
2018-10-11T08:21:08-0400


Answer:

V=540 ml = 0.54 L

Cm= 0.1115 M

m=9.987 g (mass of residue)

m(H2O) = m(mass of residue) – m(mass of CaCl2)

Formula:

n=m/Mr,. Cm=n/V. Therefore Cm=m/Mr*V ,. m=Cm*Mr*V

m=0.1115 x 111 x 0.54

m= 6.68331 g ( mass of CaCl2) (0.06 moll)

m= 9.987 – 6.68331 = 3.30369 g (H2O) (0.18 moll)

n1:n2 = 0.06 : 0.18 = 1:3

CaCl2 x 3H2O